OpenStudy (anonymous):

Use the inner product =f(−3)g(−3)+f(0)g(0)+f(3)g(3) in P2 to find the orthogonal projection of f(x)=2x^2+4x+2 onto the line L spanned by g(x)=3x^2−5x+8.

OpenStudy (anonymous):

If you were asked to do the same thing with vectors, do you know where you would start?

OpenStudy (anonymous):

yep..don't understand the format

OpenStudy (anonymous):

and my teacher leaves it to us to teach ourselves

OpenStudy (anonymous):

Well, think about it in terms of vectors in the following way. Let \[<1, 4, 5> \text{ and } <0, 2,0> \] be two vectors, A and B respectively. If I want the orthogonal projection of A onto B, I first turn B into a unit vector by dividing by the magnitude, so it becomes \[<0,1,0>\] Next, I take the inner product of this with the vector A, yielding \[<0,1,0> \cdot <1,4,5> = 4\] Finally I multiply this scalar by my old unit vector, yielding an orthogonal projection of \[<0,1,0> \cdot 4 = <0,4,0>\] Do you follow this?

OpenStudy (anonymous):

yeah i do, any idea on how to apply it to this question?

OpenStudy (anonymous):

Yes. Do exactly the same thing. You must abstract away from Euclidean space, but the concepts are the same. You are given the definition of some inner product above. It's not a dot product, like you have with vectors, but it's a perfectly valid inner product. So working through it just like we did with the vectors:

OpenStudy (anonymous):

I have two objects, f and g. I want the orthogonal projection of f onto g. First, I make g a unit object by dividing by its magnitude (what do we mean by magnitude? Simply the object divided by the square root of its inner product with itself). Next, I take the inner product of this object with my other object f. Finally, I multiply this scalar by the unit object g.

OpenStudy (anonymous):

Ill work the problem out on here, and let me know if you have a problem with it. ill show you what iv done. my hw is online so it gives instant feedback

OpenStudy (anonymous):

if you dont mind

OpenStudy (anonymous):

((<8,2,32><50,8,20>)/sqrt(2964))dot <50,8,20>

OpenStudy (anonymous):

thats essentially what i did after just plugging in f(-3),f(0),f(3),etc

OpenStudy (anonymous):

i come up with 19.3966...which is incorrect

OpenStudy (anonymous):

Sure, I'll work it out here anyway. The magnitude of g is <g,g> which is apparently \[|g| = \sqrt{<g,g>} = \sqrt{g(-3)^2 + g(0)^2 + g(3)^2} = \sqrt{50^2 + 8^2 + 20^2}\approx 54.44\] so I define the "unit function" \[g_u = \frac{g}{|g|} \] and take the inner product of this with f. That is, \[<f,g_u> = \frac{f(-3)g(-3) + f(0)g(0) + f(3)g(3)}{|g|}= \frac{400 + 16 + 640}{54.44} \approx 19.4 \] Finally, I multiply this scalar by my unit object to find my projection P: \[P = 19.4\cdot g_u = 58.19x^2 - 96.99 x + 155.18\]

OpenStudy (anonymous):

Oops. I meant to say "the magnitude of g is sqrt(<g,g>)... " the calculation is right though. Unless I made a typo :)

OpenStudy (anonymous):

nope pellet. hahah I did

OpenStudy (anonymous):

Just a sec

OpenStudy (anonymous):

on my web homework they are looking for a single magnitude....lol thanks for the explanation. i might have to ask my teacher about it, not seeing what i'm doing incorrect